### v1

Abstract
Beam monitoring is crucial for particle accelerators to achieve high luminosity. We describe an upgrade of the LABM, a beam monitoring device utilizing observation of the beamstrahlung, radiation emitted by a beam of charged particles when it accelerates in the electromagnetic field of another beam of charged particles.
I Introduction
The main task of the modern particle physics is a search for the phenomena beyond the Standard Model (BSM). Since BSM effects are expected to be small, we need large data samples, which requires a stable accelerator operation with high luminosity. To achieve it the beam monitoring is crucial.
SuperKEKB [1, 2] is an intersecting double storage ring particle accelerator with a circumference of 3.016 km that collides electrons and positrons to provide luminosity for the Belle II experiment [3]. SuperKEKB is the successor of KEKB [4]. While KEKB achieved a maximum peak luminosity of $2.11\times 10^{34}\,\textrm{cm}^{-2}s^{-1}$ [4], SuperKEKB will attempt to reach a peak luminosity of $8\times 10^{35}\,\textrm{cm}^{-2}s^{-1}$ [1], about 40 times larger than its predecessor. Since the luminosity is strongly dependent on beam-optical parameters at the IP, it is important to have direct measurement of such parameters. Such information is especially crucial to a nano-beam collider as SuperKEKB, where there is strong sensitivity to small parameter changes, in order to maximize the luminosity extracted for bunch crossing. The Large Angle Beamstrahlung Monitor (LABM) is a device designed to measure the beamstrahlung emitted at the Interaction Point (IP) of an $e^{+}e^{-}$ collider. Beamstrahlung is the radiation emitted by two beams of charged particles due to their electromagnetic interaction [5]. Beamstrahlung is directly related to the size and configuration of the beams, and provides direct information on the beams at the IP. Specifically, vertically unequal beams create an excess of y-polarized light as seen by the telescope observing the larger beam (conversely, y-polarized light from the smaller beam will decrease).
II LABM at SuperKEKB
The device that we want to describe in this paper is the Large Angle Beamstrahlung Monitor (LABM). Beamstrahlung is the radiation emitted by a beam of charged particles when it accelerates in the electromagnetic field of another beam of charged particles at the interaction point (IP) [5]. Beamstrahlung polarization and spectra are closely related to the beam configuration at the IP, and therefore the LABM can be used to passively monitor the position and size of the beams directly at the IP. The beamstrahlung was first observed at SLC [6], the linear electron–positron collider at SLAC. The first LABM prototype was successfully tested at CESR [7], a $e^{+}e^{-}$ collider located at Cornell University, and operated between 2007 and 2008. An upgraded version of the LABM was installed at SuperKEKB in 2015 [8] and operates since 2016 [9].
LABM operating at SuperKEKB accelerator consists of four telescopes pointed at the Belle II IP, at around 8.5 mrad observation angle from the beam axis. On each side of Belle II detector there are two telescopes on the top (azimuthal position of 90 degrees) and bottom (270 degrees) of one of the beam pipes. This redundancy improves the systematics and allows to be sensitive to vertical asymmetries of the particle beams.
Beamstrahlung is intercepted inside the beam pipe with vacuum mirrors made of beryllium at about 5 meters downstream from the IP and extracted through a special window, as shown in Fig. 1. The size of the mirrors is $\times2$.8 mm2, which corresponds to a square angular acceptance once the mirror inclination is taken into account. Then the light travels inside a series of straight aluminum pipes with mirrors at pipe joints until it arrives to an optical box located outside of the interaction region. There are two optical boxes, one for the electron beam (“Oho” side of the detector) and one for the positron beam (“Nikko” side of the detector). The first mirror after the vacuum window, referred to as a “primary mirror”, is movable by two stepper motors and may be controlled remotely, which allows to perform an angle scan in search of the signal.
Figure 1: Two vacuum mirrors are used to extract the beamstrahlung light from the vacuum beam pipe through small windows. The beamstrahlung light coming from the IP is indicated with arrows. The light is then guided towards the detector by a system of pipes and mirrors, which is not shown here.
In the optical boxes the light is split into two transverse polarization components. The beamstrahlung is expected to be highly polarized, so that we can observe the signal in one polarization and use the other for the background estimation. In the classical LABM design used until 2023 each component was observed at four different average wavelengths by photomultiplier tubes (PMT). See Refs. [8] and [9] for more details of that design. After an upgrade in 2023, which is described in this paper, the PMTs were replaced with cameras.
After several years of studies and upgrades in 2022 we could successfully demonstrate that the LABM measurements of beamstrahlung signal are of good quality and are correlated with key beam parameters. Due to the problems with getting the proper IP spot for the positron beam, only two telescopes monitoring the electron beam were used, i.e. measurements from 16 PMTs. Data sample was collected with telescopes looking at the IP without moving for 11 days. Only Physics run data with currents above 100100\,mA was used. The results were obtained with a Machine Learning (ML) using a Neural Network (NN). External measurements of the beams sizes from the accelerator group and luminosity from the detector group were used for training. The mean absolute error (MAE) of prediction of the beam parameters on a data sample not used in the training was at a few percent level [10].
III The 2023 upgrade
The classic LABM design had several drawbacks. First, PMTs provide only a point-like measurement. To get an image of the IP including areas surrounding the spot a scan was necessary which could take significant time of the order of hours, while beam conditions change every second. It was possible to take measurements in one point without moving but it provided a limited information. Besides, over longer periods the spot might move and tracking it needed a new scan.
Second, since the optical boxes are located outside of the beamline, it is necessary to use long optic paths with many mirrors which need adjustment. The adjustment procedure is error prone and there is always a risk of choosing instead of a true spot a reflection, a fake spot etc.
In 2023 we performed a major LABM upgrade and replaced PMTs with CMOS a2A1920-51gcPRO cameras from Basler [11]. This allowed us to capture the image of the IP spot and backgrounds together at once. Because of this and since the cameras make an image (measurement) in three ranges of the spectrum at once, most of the internals of the optic boxes became unnecessary.
We expect the spot to have a size of approx. $4\times 4\,$mm. We would also like to have an image of the area around the spot to estimate backgrounds. Since the size of the sensor is $6.6\times 4.1\,$mm, we use a focusing lens after the Wollaston prism. The focal length of the lens is 4040\,cm, and the cameras are installed at 3030\,cm from the prism, as shown in Fig. 2. Gratings, mirrors, PMTs, electronics and conveyor belts were removed from the optical boxes.
We also replaced primary mirrors drives: stepper motors, controllers and cables to avoid interference problems which we observed before. The optic path remained the same for this data taking period.
1233
Figure 2: Optic box with cameras. The light enters the box as indicated by an arrow, the Wollaston prism (1) splits the light polarizations, the lens (2) focuses the light on the cameras (3).
Before taking data we performed angle scans in search for the light from the IP, which we call a “spot”. The results of the scans are shown in Fig. 3. For every position of the stepper motors ($x$- and $y$-axes on the plots) a sum of all pixels of an image was calculated. This value is shown for both cameras of each telescope as PMT1 and PMT2. After a rough scan in a wide angle range we performed a fine scan around a found spot. Due to the hysteresis effect in the primary mirrors’ drives the exact position of the spot may change a little from a scan to a scan, thus final positioning was performed by hand to fit a spot into a camera sensor. Examples of the spot images for all eight cameras are shown in Figs. 4 and 5. Each image is a combination of three colors: red, green and blue. As can be seen the position of the spot is slightly different for different colors. This happens since Wollaston prism is chromatic and the separation angle between the ordinary and extraordinary rays depend on the wave length. When a primary mirror moves the location of the spot changes for two cameras capturing two light polarizations simultaneously. This and the Wollaston prism chromaticity limits our ability for the ideal spot positioning.
Data used in this study was taken in May – June 2024. In total seven good runs were taken with the total duration of 23 days. All data with currents above 100 mA were considered for analysis.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: The results of the scans for x- (PMT1) and y- (PMT2) polarizations for Oho down (upper left), Oho up (upper right), Nikko up (lower left), Nikko down (lower right).
Refer to caption Refer to caption
Refer to caption Refer to caption
RGBRGBRGBRGB
Figure 4: Examples of images of the spot for three color components (R, G and B) for Oho down x-polarization (upper left), Oho down y-polarization (upper right), Oho up x-polarization (lower left), Oho up y-polarization (lower right).
Refer to caption Refer to caption
Refer to caption Refer to caption
RGBRGBRGBRGB
Figure 5: Examples of images of the spot for three color components (R, G and B) for Nikko down x-polarization (upper left), Nikko down y-polarization (upper right), Nikko up x-polarization (lower left), Nikko up y-polarization (lower right).
IV ML model description
To obtain the results we again used an ML approach. There were two methods to process collected images. The first one featured a convolutional neural network (CNN) which used as input scaled to $240\times 192$ pixels but otherwise unprocessed images, denoted as “full images” (FI). The second method featured a separate library openCV [12], which was used to describe each image with an array of 768 parameters, denoted as a “reduced image” (RI). Then this array was fed to a NN. For most variables the FI method gives better results than the RI one, but it requires much more computational resources and time for training.
We tried many different models for the FI method. We still continue this process. The default model for now consists of three convolutional layers, one fully connected layer and an output layer. The total number of neurons in this model is approx 11M. The RI method uses a model with a single hidden fully connected layer. The total number of neurons in this model is approx 665K. The models were implemented using Keras [13], a high level abstraction library that works on top of the low level TensorFlow [14] compute engine.
The data set used consists of about 130000 LABM and SuperKEKB measurements, which are split in about 90000 points for training, 25000 for validation, and 13000 for testing of the model, obtained over 23 days every eight seconds. These measurements are randomly shuffled before being used, meaning that there is no time correlation between successive points in the data set.
V Results
In this section we show the results of the NN models and compare these results with those provided by the traditional Linear Regression (LR):
$y=\beta_{0}+\sum_{i}\beta_{i}x_{i}\;,$ (1)
where independent variables $x_{i}$ correspond to the number of pixels of scaled images in FI method and to the number of parameters in an image description in RI method. A LR approach is feasible as, historically, we have always seen high linearity with beam current and other beam parameters in zero beamstrahlung conditions (that is, when only one beam was present in the accelerator). We also compare our results with those obtained by the previous LABM configuration from [10].
In this study we try to reproduce the following accelerator-related variables: luminosity, currents, beam sizes and emittances.
V.1 Specific luminosity
Although luminosity is not a beam parameter, it is strictly related to the beam parameters at the IP and it constitutes, together with the energy, one of the two figures of merit of a collider. Therefore, it will be the first experimental measurement that we will try to reproduce with our models. The absolute luminosity at SuperKEKB is measured by the Electromagnetic Calorimeter (ECL) monitor [15], located in the Belle II detector. The ECL measures Bhabha events and, following calibration, provides an absolute value for the luminosity. For the direct comparison with the previous configuration we will use the specific luminosity, defined as:
$\mathcal{L}_{s\!p}=\frac{f_{0}}{2\pi\Sigma_{x}\Sigma_{y}}$ (2)
where $f_{0}$ (0.1 MHz for SuperKEKB) is equal to the single bunch revolution frequency, and $\Sigma_{i}$ $(i=x,y)$ are the convoluted beam sizes, corresponding to the quadrature sum of the two beam sizes at the IP: $\Sigma_{i}^{2}=\sigma_{i,1}^{2}+\sigma_{i,2}^{2}$. It is obtained from the regular formula of the luminosity dividing by the factor $N_{b}N_{1}N_{2}$, i.e. the number of bunches times the product of the numbers of particles per bunch of the two beams. In this way, the specific luminosity is independent from the beam currents and from the number of bunches present in the rings. Fig. 6 shows the specific luminosity as predicted by LR.
In the previous publications it was shown that in the SuperKEKB beam crossing situation the specific luminosity should closely track the variable
$\sigma_{y,e\!f\!f}=\frac{\sigma_{y,1}\sigma_{y,2}}{\Sigma_{y}}.$ (3)
It is shown in Fig. 7, where the data points are sorted in the ascending order of the variable $\sigma_{y,e\!f\!f}$. The MAE for the plots is 2.2%, 1.16% and 1.3%, respectively, which shows clear improvement with the new setup.
Refer to caption Refer to caption Refer to caption
preliminarypreliminary
Figure 6: Specific luminosity obtained with the previous setup [10] (left), and with the upgraded setup with full images CNN (center) and reduced images NN (right).
Refer to caption Refer to caption Refer to caption
preliminarypreliminary
Figure 7: $\sigma_{y,e\!f\!f}$ obtained with the previous setup [10] (left), and with the upgraded setup with full images CNN (center) and reduced images NN (right).
V.2 Vertical beam size
The beam size at SuperKEKB is measured by the X-ray monitor (XRM) [16]. The XRMs are installed in both the HER ($e^{-}$) and LER ($e^{+}$) rings, at 641.4 and 1397.7 m respectively from the IP. Using the Twiss parameters at their location, it is possible to obtain an estimate of the emittance, and through the optical transfer matrices it is possible to estimate the beam size at the IP, which is the quantity we are interested in.
Figs. 8 and 9 show the results for the vertical sizes of the beams $\sigma_{y}$. We are also interested in the LER/HER ratio on $\sigma_{y}$ at IP. In fact, KEKB had a vertical beam size for the LER that was consistently larger than the corresponding one for the HER. This corresponds to one beam being unfocused, causing significant luminosity degradation, which is an effect that we want to prevent at SuperKEKB. Fig. 10 shows the ratio $\sigma_{y}$ LER/HER. The plots related to horizontal beam widths and their ratio are shown in Figs. 11 to 13.
Refer to caption Refer to caption Refer to caption
preliminarypreliminary
Figure 8: LER $\sigma_{y}$ obtained with the previous setup [10] (left), and with the upgraded setup with full images CNN (center) and reduced images NN (right).
Refer to caption Refer to caption Refer to caption
preliminarypreliminary
Figure 9: HER $\sigma_{y}$ obtained with the previous setup [10] (left), and with the upgraded setup with full images CNN (center) and reduced images NN (right).
Refer to caption Refer to caption Refer to caption
preliminarypreliminary
Figure 10: LER/HER ratio $\sigma_{y}$ obtained with the previous setup [10] (left), and with the upgraded setup with full images CNN (center) and reduced images NN (right).
Refer to caption Refer to caption Refer to caption
preliminarypreliminary
Figure 11: LER $\sigma_{x}$ obtained with the previous setup [10] (left), and with the upgraded setup with full images CNN (center) and reduced images NN (right).
Refer to caption Refer to caption Refer to caption
preliminarypreliminary
Figure 12: HER $\sigma_{x}$ obtained with the previous setup [10] (left), and with the upgraded setup with full images CNN (center) and reduced images NN (right).
Refer to caption Refer to caption Refer to caption
preliminarypreliminary
Figure 13: LER/HER ratio $\sigma_{x}$ obtained with the previous setup [10] (left), and with the upgraded setup with full images CNN (center) and reduced images NN (right).
Table 1: Comparison of the results
Model PMT NN RI NN FI NN
$\mathcal{L}_{spec}$ 3.4% 1.6% 1.3%
$\sigma_{y,e\!f\!f}$ 2.2% 1.3% 1.16%
$\sigma_{x,LER}$ 0.7% 1.2% 1.13%
$\sigma_{x,HER}$ 0.9% 0.6% 0.6%
$\sigma_{x,LER}/\sigma_{x,HER}$ 1.1% 1.4% 1.3%
$\sigma_{y,LER}$ 3.7% 2.4% 2.0%
$\sigma_{y,HER}$ 3.4% 2.6% 1.8%
$\sigma_{y,LER}/\sigma_{y,HER}$ 4.8% 4.3% 3.1%
VI Conclusion
VII Future developments
Using cameras instead of PMTs opens new possibilities for the LABM development. The optical box may be miniaturized and moved next to the beamline, in which case we greatly reduce number of mirrors in the telescopes. It makes the device adjustment easier, number of reflections and false spots smaller and the signal larger. A possible disadvantage of this change is the radiation level the device will have to cope with. However our estimations show that the cameras should be able to survive it for at least several years.
With a minimal exposition of the cameras of order of 10 ms, we can measure very fast beam changes, though we may be limited by the amount of light and the sensor sensitivity.
Another issue is how we get the desirable results from the beamstrahlung observation. At the moment LABM relies on the external measurements from the accelerator group and Belle II detector for the training of the NNs. Another direction of the device development is to make completely independent measurements of the beam parameters by using beamstrahlung simulation. To realize this idea we communicate with two groups working on beam simulation packages, SAD [17] and Xsuite [18]. At the moment none of them provides complete photon simulation but both may add it in the future.
EIC is the electron-ion collider at BNL. Even though only one beam is electron and LABM usage for the proton beam is yet be studied, we showed that even monitoring a single beam may give valuable information. In future LABM-like device may be installed on EIC [19].
Another accelerator which may greatly benefit from the beamstrahlung monitoring is FCC-ee [20]. The intensity of the beam and therefore beamstrahlung brightness as well as radiation levels around the accelerator make it challenging, but the benefits of the beam monitoring through beamstrahlung are indisputable.